# Automated Proof Generation for Rust Code via Self-Evolution

### v1

@@ D.3.1 SAFE’s Advantages Compared to GPT-4o
D.3.2 The Effectiveness of Our Self-Debugging Mechanism
@@
Listing 10 shows the initial proof generated by SAFE, which includes loop invariants: (1) the index does not exceed the length of the given array; (2) all elements before the current index satisfy the ‘‘sorted’’ property. However, during each iteration of this loop, this program checks the current element against the next one and terminates the loop when index equals arr.len() - 2. As a result, the loop invariants fail to account for the last element, leading to the error described in Lines 29-41.
@@
D.3.3 Limitations
We reported that SAFE failed to generate correct proof for 20.86% of the programs in VerusBench, even after sampling 100 proofs and 100 debugged proofs for each incorrect proof (i.e., Accuracy@100). In Listing 12 and Listing 13, we analyze an example where SAFE fails to generate a correct proof. This program removes odd numbers from a given array by scanning the input array arr and appending the even numbers to a new array, even_list.
@@ D.3.3 Limitations
The root cause of this error is that proving this program needs to add assert statements and leverage the reveal feature of Verus (Doc, 2024). We suspect the reveal feature was never synthesized in any finetuning dataset, so the final SAFE model also lacks sufficient domain knowledge to prove this program. This limitation can be substantially mitigated by incorporating a larger and more diverse training set during the initial round of self-evolution.
@@
D.3.4 An Example Proof with a Low-Quality Specification
In Section 4.2.4, we demonstrate that the quality of specifications contribute substantially to the end-to-end effectiveness of SAFE. Listing 14 presents an example proof with a low-quality specification. This program computes the greatest common divisor of two integer number using the Euclidean Algorithm. However, this specification merely requires that the result be greater than zero, ‘‘result > 0’’, without ensuring that the result should be divisors of both inputs a and b.
Programs with such low-quality specifications might lead to extremely simple proofs, which are very different from proofs in more realistic tasks. For example, the proof annotations synthesized in Lines 16-17 simply copy the function pre-condition — a > 0 and b >= 0. For more realistic tasks, proofs can rarely get done this way. Consequently, putting many proofs with this quality (i.e., all resulted from low-quality specifications) into the fine-tuning dataset does not help model’s evolution.

### v2

@@ F.1 SAFE’s Advantages Compared to GPT-4o
F.2 The Effectiveness of Our Self-Debugging Mechanism
@@
Listing 10 shows the initial proof generated by SAFE, including loop invariants: (1) the index does not exceed the length of the given array; (2) all elements before the current index satisfy the ‘‘sorted’’ property. However, during each iteration of this loop, this program checks the current element against the next one and terminates the loop when index equals arr.len() - 2. As a result, the loop invariants fail to account for the last element, leading to the error described in Lines 29-41.
@@
F.3 An Ineffective Example
We report that SAFE fails to generate correct proof for 20.86% of the programs in VerusBench, even after sampling 100 proofs and 100 debugged proofs for each incorrect proof (i.e., Accuracy@100). In Listing 12 and Listing 13, we analyze an example where SAFE fails to generate a correct proof. This program removes odd numbers from a given array by scanning the input array arr and appending the even numbers to a new array, even_list.
@@ F.3 An Ineffective Example
The root cause of this error is that proving this program needs to add assert statements and leverage the reveal feature of Verus (Doc, 2024). We suspect that the reveal feature is never synthesized in any finetuning dataset, so the final SAFE model also lacks sufficient domain knowledge to prove this program. This limitation can be substantially mitigated by incorporating a larger and more diverse training set during the initial round of self-evolution.
@@
F.4 An Example Proof with a Low-Quality Specification
In Section 4.3.3, we demonstrate that the quality of specifications contributes substantially to the end-to-end effectiveness of SAFE. Listing 14 presents an example proof with a low-quality specification. This program computes the greatest common divisor of two integer numbers using the Euclidean Algorithm. However, this specification merely requires that the result be greater than zero, ‘‘result > 0’’, without ensuring that the result should be divisors of both inputs a and b.
Programs with such low-quality specifications might lead to extremely simple proofs, which are very different from proofs in more realistic tasks. For example, the proof annotations synthesized in Lines 16-17 simply copy the function precondition—a > 0 and b >= 0. For more realistic tasks, proofs can rarely get done this way. Consequently, putting many proofs with this quality (i.e., all resulted from low-quality specifications) into the fine-tuning dataset does not help a model’s evolution.